Circular No. 7088

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1999C IN ANONYMOUS GALAXY A. Gal-Yam and D. Maoz, Tel Aviv University, report for the Wise

Observatory Optical Transients Search (see IAUC 7055) their discovery of an apparent supernova on unfiltered CCD images taken with the Wise Observatory 1-m telescope on Jan. 14. SN 1999C (R = 19.6) is located at α = 10^{h}08^{m}51^{s}.21, δ = +71^{o}10′42′′.9 (equinox 2000.0), which is in the field of the galaxy cluster Abell 914 (z = 0.195). The object is located 0′′.9 east of, and at the same δ as, the host galaxy’s center. SN 1999C was invisible in images obtained on 1998 Mar. 31 (limiting mag R ∼ 22.0). The object was already detectable in images obtained on 1998 Dec. 23 (R = 19.8) and was confirmed on CCD images taken by A. Gal-Yam on 1999 Jan. 15 with the same telescope. Finding charts can be obtained via ftp://wise-

obs.tau.ac.il/pub/avishay/SN/sn1999#.

SUPERNOVA 1999D IN NGC 3690 Y. L. Qiu, Q. Y. Qiao, and J. Y. Hu, Beijing Astronomical Observatory

(BAO), report their discovery of an apparent supernova in the course of the BAO Supernova Survey with the 0.6-m telescope at Xinglong Station. The new star is located at α = 11^{h}28^{m}29^{s}.50, δ = +58^{o}33′40′′.6 (equinox 2000.0), which is ∼ 14′′.1 west and 12′′.6 north of the bright knot of the western part of NGC 3690 (cf. IAUC 6859; note that SN 1998T was near the center of NGC 3690). The object was found and confirmed on unfiltered CCD images taken on Jan. 16.83 and 17.65 UT that show magnitudes of 15.6 and 15.3, respectively. CCD frames taken on 1998 Dec. 30 show no star at the position of the supernova (limiting mag ∼ 18.0). W.-d. Li, University of California at Berkeley, reports that KAIT images show the presence of SN 1999D already on 1999 Jan. 11.

SUPERNOVA 1999B IN UGC 7189 S. Jha, P. Garnavich, P. Challis, and R. Kirshner, Center for Astro-

physics, report that a spectrum of SN 1999B (cf. IAUC 7086), obtained on Jan. 16.5 UT by M. Calkins with the F. L. Whipple Observatory’s 1.5-m telescope (+ FAST spectrograph), reveals it to be a type-II supernova. The spectrum exhibits a blue continuum with P-Cyg Balmer lines and a weak He I (rest wavelength 587.6 nm) absorption. Adopting the NASA/IPAC Extragalactic Database recession velocity for UGC 7189 of 1683 km/s implies a photospheric expansion velocity of 6700 km/s for the supernova, measured to the trough of the Hα line profile.

1999 January 18  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1999 CBAT  Daniel W. E. Green